OpenStudy (anonymous):

ok say i have xsquared minus x -12 how would i factor the quadratic equation?

OpenStudy (anonymous):

fully steps

OpenStudy (anonymous):

determine the prime factors of the constant term from those try to construct the coefficient of x using either addiition of some terms or subtraction of multiples of obtained factors then re write the quadratic take common terms outside now you have the factors for example - 12 can have factors \[\pm1,12,2,3,4,6\] from these i can choose -4 and +3 to create -1 now the quadartic becomes x^2-4x+3x-12 x(x-4)+3(x-4) =(x-4)(x+3) this is how i do it may sound bit crazy but i believe most people do the same

OpenStudy (anonymous):

I always think.. \[Ax^2+Bx+C\] what added together makes B and when times together makes C.. not as nice as Vicky007's answer..but easy to remember.

OpenStudy (anonymous):

only way to factor this is as \[(x-\sqrt{12})(x+\sqrt{12})\]as the difference of two squares. halas

OpenStudy (anonymous):

its only an expression though shouldnt it be muchh easier

OpenStudy (anonymous):

easier than what? are you trying to solve \[x^2-12=0\] or are you trying to factor \[x^2-12\] i am not sure which. but you will notice that there is no "x" term here as compared to say \[x^2-x-12\] so it is simple to factor

OpenStudy (amistre64):

x^2-(x-12) is what i read it

OpenStudy (amistre64):

"xsquared minus x -12"

OpenStudy (anonymous):

oooooooooooooooooooh i should quit while i am behind!

OpenStudy (anonymous):

its just xsquared minus x minus 12

OpenStudy (anonymous):

\[x^2-x-12\] like that?

OpenStudy (anonymous):

all i wanna know is the easiestway to solve it

OpenStudy (amistre64):

dont use english; use the math notation

OpenStudy (anonymous):

i read \[x^2-12\] doh

OpenStudy (amistre64):

there is no mathematical "easiest" way to solve it

OpenStudy (anonymous):

there is no easy way other than thinking.

OpenStudy (amistre64):

there are a variety of methods; but math doesnt pick one over the other

OpenStudy (anonymous):

two numbers whose product is -12. one plus one minus

OpenStudy (anonymous):

sum is -1

OpenStudy (anonymous):

ok whats a method that is easilhy understood

OpenStudy (anonymous):

ok rather say i have xsqua

OpenStudy (anonymous):

if you want a sure fire method complete the square and then you don't have to think

OpenStudy (amistre64):

one that will always produce and answer is the quadratic formula: (-b +- sqrt(b^2 -4ac))/2a

OpenStudy (amistre64):

\[ax^2+bx+c=0\] \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

ok then after i plug all that in thats my answer

OpenStudy (anonymous):

what does the - b + or minus mean

OpenStudy (amistre64):

yep; the formula is derived thru alot of steps; but in the end, that is what it amounts to

OpenStudy (amistre64):

plug in the numbers and hit enter on a calculator :)

OpenStudy (amistre64):

when we have 2 exact expression that differ only by an operation; it is shorthand notation to just include them as one sign

OpenStudy (amistre64):

x+2 and x-2 , can be written as

OpenStudy (amistre64):

x +- 2

OpenStudy (anonymous):

indeed\

OpenStudy (amistre64):

a quadratic equation will produce at most; 2 answers, that will differ by a + or - operation right there

OpenStudy (anonymous):

ok so from xsquared minus x-12 i got x=-6.5

OpenStudy (anonymous):

well this is one of the longest reply thread i have ever seen

OpenStudy (anonymous):

my answer always comes out as a relally long decimal